# AceReason-Nemotron: Advancing Math and Code Reasoning through Reinforcement Learning

### v1

@@ 3.1 Framework
We adopt the GRPO algorithm (Shao et al., 2024), as used in DeepSeek-R1, instead of PPO (Schulman et al., 2017), due to its simplicity and the advantage of not requiring a separate value function model. For each question-answer pair $(q,a)$, we sample the policy model $\pi_{\theta_{\text{old}}}$ to generate a group of $G$ individual rollouts $\{o_{i}\}_{i=1}^{G}$. We assign a reward score $S_{i}=S(o_{i},a)$ to each response $o_{i}$, given the oracle answer $a$, using a rule-based reward function $S$, and compute the advantage as the normalized reward scores across the group $\{S_{i}\}_{i=1}^{G}$.
@@
where
$r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})},\leavevmode\nobreak\ \hat{A}_{i,t}=\frac{S_{i}-\text{mean}(\{S_{i}\}_{i=1}^{G})}{\text{std}(\{S_{i}\}_{i=1}^{G})}.$ (2)
Our experiments strictly follow the on-policy training method, performing only one gradient update after the model generates a group of $G$ rollouts, to ensure stable RL training and prevent entropy collapse. We also remove the KL divergence term by setting $\beta=0$.
We started RL experiments from distilled reasoning models such as DeepSeek-R1-Qwen-Distilled-7B/14B (Guo et al., 2025), ensuring that the experiments were conducted under controlled conditions without introducing variations in distillation data or fine-tuning compute. We use RL framework veRL (Sheng et al., 2024) with implementation of token-level loss (Yu et al., 2025) of GRPO and inference engine vLLM (v0.7.3) (Kwon et al., 2023). Our custom modification includes adding math and code reward functions (verifiers) to the implementation. All experiments are conducted using 128 NVIDIA H100 GPUs.
@@ 4.3.7 Does RL Improve pass@k or pass@1 over the SFT Model?
Recent studies (Shao et al., 2024; Yue et al., 2025) suggest that RL primarily improves pass@1 accuray over SFT model ((DeepSeek-R1-Distilled) without significantly impacting pass@$k$. However, Figure 6 demonstrates that RL consistently improves pass@$k$ score (from $k=8$ to $k=64$) for both 7B and 14B models on LiveCodeBench v5 and v6, with an improved pass@k scores maintaining a $10\%$ margin from pass@$8$ to pass@$64$. On AIME24/25, we found the 14B model also shows better pass@k, although the AIME25 gap on 7B is closed by the SFT model at pass@$64$.

### v2

@@ 3.1 Framework
We adopt the GRPO algorithm (Shao et al., 2024), as used in DeepSeek-R1, instead of PPO (Schulman et al., 2017), due to its simplicity and the advantage of not requiring a separate value function model. For each question-answer pair $(q,a)$, we sample from policy model $\pi_{\theta_{\text{old}}}$ to generate a group of $G$ individual rollouts $\{o_{i}\}_{i=1}^{G}$. We assign a reward score $S_{i}=S(o_{i},a)$ to each response $o_{i}$, given the oracle answer $a$, using a rule-based reward function $S$. We employ the token-level policy gradient loss variant of GRPO, as introduced by Yu et al. (2025),
@@
therein,
$r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})},\leavevmode\nobreak\ \hat{A}_{i,t}=\frac{S_{i}-\text{mean}(\{S_{i}\}_{i=1}^{G})}{\text{std}(\{S_{i}\}_{i=1}^{G})}\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \forall t,$ (2)
where $r_{i,t}(\theta)$ is the token-level importance weight, and the token-level advantage $\hat{A}_{i,t}$ within each response is uniformly assigned with the value of normalized reward score across the group $\{S_{i}\}_{i=1}^{G}$.
Our experiments strictly adhere to on-policy training by performing a single gradient update after generating a group of $G$ rollouts. This approach ensures stable RL training and helps prevent entropy collapse. Consequently, the policy used for data collection matches the current policy, i.e., $\pi_{\theta_{\text{old}}}(\cdot\mid q)=\pi_{\theta}(\cdot\mid q)$, and the importance weight $r_{i,t}(\theta)=1$. Additionally, we eliminate the KL divergence term by setting $\beta=0$, thus the GRPO objective becames REINFORCE objective (Williams, 1992) with group-normalized rewards,
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{(q,a)\sim\mathcal{D},\leavevmode\nobreak\ \{o_{i}\}_{i=1}^{G}\sim\pi_{\theta}(\cdot\mid q)}\Bigg{[}\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\hat{A}_{i,t}\Bigg{]}.$ (3)
So, the update rule becomes,
$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{(q,a)\sim\mathcal{D},\leavevmode\nobreak\ \{o_{i}\}_{i=1}^{G}\sim\pi_{\theta}(\cdot\mid q)}\Bigg{[}\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\cdot\hat{A}_{i,t}\Bigg{]}.$ (4)
We started RL experiments from distilled reasoning models DeepSeek-R1-Qwen-Distilled-7B/14B (Guo et al., 2025), ensuring that the experiments were conducted under controlled conditions without introducing variations in distillation data or fine-tuning compute. We use the veRL framework (Sheng et al., 2024), which implements token-level loss for GRPO, and employ the vLLM inference engine (v0.7.3) (Kwon et al., 2023) for sample generation. Our custom modification includes adding math and code reward functions (verifiers) to the implementation. All experiments are conducted using 128 NVIDIA H100 GPUs.
@@ 3.3.2 Training Process
3.4 Summary of Training Curriculum
We use DeepSeek-R1-Distill-Qwen2.5-7B and 14B as our initial SFT models. To integrate math-only and code-only RL, we first perform math-only RL with stage-wise length extension from 8K to 24K. Next, we apply code-only RL, further extending the length from 24K to 32K. Finally, we conduct math-only RL with a response length of 32K. We find that this training curriculum is slightly more effective and efficient in practice than first performing math-only RL from 8K to 32K, followed by code-only RL from 24K to 32K.
@@ 4 Evaluation
We provide the experimental setup, results, and extensive analysis in this section. The main results are summarized in Table 2.
@@ 4.3.7 Does RL Improve pass@k or pass@1 over the SFT Model?
Recent studies (Shao et al., 2024; Yue et al., 2025) suggest that RL primarily improves pass@1 accuray over SFT model (DeepSeek-R1-Distilled) without significantly impacting pass@$k$. However, Figure 6 demonstrates that RL consistently improves pass@$k$ score (from $k=8$ to $k=64$) for both 7B and 14B models on LiveCodeBench v5 and v6, with an improved pass@k scores maintaining a $10\%$ margin from pass@$8$ to pass@$64$. On AIME24/25, we found the 7B and 14B model also show better pass@k compared to the SFT model across all $k$. For both AIME and LiveCodeBench, we generated 128 responses for each question and randomly sampled $k$ to calculate the pass@$k$ result with an average of 100 runs to reduce variance.